Find the value of k such that the graphs of x + 7y = 70 and y + 3 = kx are perpendicular.
Answer by stanbon(75887)

Find the value of k such that the graphs of x + 7y = 70 and y + 3 = kx are perpendicular.
----------------------
Solve each equation for "y":
y = (-1/7)x + 10
---
y = kx - 3
---
If the 2nd equation is perpendicular to the 1st,
(-1/7)k = -1
(1/7)k = 1
k = 7
------------------------
Cheers,
Stan H.